# ../Bibtex/sources.bib

### v1

@@ 1. Introduction
The Oguiso-Yu K3 surface $X_{OY}$ is the elliptic K3 surface whose weierstrass equation is
@@ Acknowledgements
The first named author would like to thank Janko Böhm for sharing his insights. We would further like to thank Anne Frühbis-Krüger for collaboration and numerous helpful discussions on the schemes package in OSCAR, as well as Max Horn, Claus Fieker, and Tommy Hofmann for their enduring support in integrating our work into the OSCAR codebase. Finally, thanks go to Hans Schönemann for resolving various issues encountered with the Singular backend in OSCAR and again to Tommy Hofmann for providing numerous improvements on number-theoretic topics throughout the preparations of this work. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 286237555 – TRR 195.
@@ 2.1. K3 surfaces
Since K3 surfaces are simply connected, linear, algebraic, and numerical equivalence agree and $\mathrm{Pic}(X)=\mathrm{NS}(X)=\mathrm{Num}(X)$. From the exponential sequence, we thus have an embedding $c_{1}\colon\mathrm{Pic}(X)\to H^{2}(X,\mathbb{Z})$ whose image is the subgroup of integral $(1,1)$-classes $H^{1,1}(X)\cap H^{2}(X,\mathbb{Z})$ by Lefschetz’ $(1,1)$-theorem. The smallest primitive sublattice of $H^{2}(X,\mathbb{Z})$ containing the period $H^{2,0}(X)$ is called the transcendental lattice $T(X)$. A K3 surface is not necessarily projective, but if it is, then $T(X)=\mathrm{NS}(X)^{\perp}$ in $H^{2}(X,\mathbb{Z})$.
@@
which preserves the intersection form and the Hodge structures, i.e. $\phi_{\mathbb{C}}(H^{2,0}(Y))=H^{0,2}(X)$. The weak Torelli theorem states that $X$ and $Y$ are isometric, if and only if $H^{2}(X,\mathbb{Z})$ and $H^{2}(Y,\mathbb{Z})$ are Hodge-isometric.
To formulate the strong Torelli theorem, we need one more definition. The Hode-isometry $\phi$ is called effective, if it satisfies one of the following three equivalent conditions:
@@
By the Riemann-Roch theorem we have that for $r\in\Delta(X)$ either $r$ or $-r$ is effective. Set $\Delta^{+}(X)=\{r\in\Delta(X)\mid r\mbox{ is effective}\}$. Then $h\in\mathcal{P}_{X}$ is ample if and only if $h.r>0$ for all $r\in\Delta^{+}(X)$. More generally, the connected components of $\mathcal{P}_{X}\setminus\bigcup\{r^{\perp}\otimes\mathbb{R}\mid r\in\Delta(X)\}$ are called Weyl-chambers. The ample cone is one such chamber. The Weyl group $W(X)\subseteq O(\mathrm{NS}(X))$ is generated by the reflections into the hyperplanes $r^{\perp}$. It acts simply-transitively on the set of Weyl-chambers.
@@ 2.2. Enriques surfaces
A complex Enriques surface $Y$ is a compact complex manifold with $h^{1}(Y,\mathcal{O}_{Y})=0$ and $b_{2}(Y)=10$. Its canonical bundle $\omega_{Y}\neq\mathcal{O}_{Y}$, but $\omega_{Y}^{\otimes 2}\cong\mathcal{O}_{Y}$. Since the canonical bundle is $2$-torsion, it induces a $2:1$ cover $X\to Y$, which turns out to be the universal covering with $X$ a K3 surface. Since $h^{1,1}(Y)=b_{2}(Y)$, it follows from Lefschetz’ Theorem on $(1,1)$-classes and the Kodaira Embedding theorem that all Enriques surfaces are projective.
@@ 3.2. The transcendental lattice
$\{f\in\mathrm{NS}(X)\mid f^{2}=0,h.f=2\}=\{e_{1},e_{2}\}.$
Since the invariant lattice is of rank $1$, $G$ must exchange $e_{1}$ and $e_{2}$ and the stabiliser of $e_{i}$ is of order $G/2=128$. Recall from [shioda] that the trivial lattice $\mathrm{Triv}(f)$ of an elliptic fibration with fiber $f\in\mathrm{NS}(X)$ is the sublattice of $\mathrm{NS}(X)$ spanned by the zero section and fiber components. The Mordell-Weil group is isomorphic to $\mathrm{NS}(X)/\mathrm{Triv}(e_{i})$. To compute the trivial lattice, one chooses any $s\in\mathrm{NS}(X)$ with $s^{2}=-2$ and $s.f=1$. Let $R$ be the root sublattice of $\langle f,s\rangle^{\perp}\subseteq\mathrm{NS}(X)$. Then the trivial lattice is isomorphic to $\mathbb{Z}\langle s,f\rangle\oplus R$. The irreducible root sublattices of $R$ give the ADE types of the singular fibers. For how to compute the fiber types on the Enriques surface $Y$ from the class of $e_{i}$ see [brandhorst-gonzalez, Prop. 5.8]. ∎
@@ 4.1. The first model
It then follows from Proposition 3.2 that $X_{OY}\cong X\cong X_{h}$ and from Proposition 3.6 that $X_{h}/j=:Y$ is isomorphic to the Oguiso-Yu Enriques surface $Y_{OY}$ of minimum entropy.
@@ 4.3. Extracting a Weierstrass model
where $\zeta_{16}$ is a primitive $16$-th root of unity. The associated elliptically fibered K3 surface $X_{e_{1}}$ is defined over $\mathbb{Q}$ and its trivial lattice over $\mathbb{Q}(\zeta_{8})$. However, in order to realize all algebraic cycles for the Néron-Severi lattice $\mathrm{NS}(X_{e_{1}})$, we still need the extension $K/\mathbb{Q}(\zeta_{8})$ of degree two generated by the algebraic element $\alpha$ with minimal polynomial given in Equation 6.

### v2

@@ 1. Introduction
Figure 1. The real points of the affine birational model for the Enriques surface sought for by Oguiso and Yu
@@
The Oguiso-Yu Enriques surface is birational to the surface with equation
$\tilde{y}^{2}=s\tilde{x}^{4}+s^{7}-s^{3}.$
Its universal cover $X_{OY}$ is the elliptic K3 surface whose weierstrass equation is
@@ Acknowledgements
Special thanks goes to Matthias Schütt for discussions and suggesting the equation for the Enriques surface after reading a first version of this paper. The first named author would like to thank Janko Böhm for sharing his insights. We would further like to thank Anne Frühbis-Krüger for collaboration and numerous helpful discussions on the schemes package in OSCAR, as well as Max Horn, Claus Fieker, and Tommy Hofmann for their enduring support in integrating our work into the OSCAR codebase. Finally, thanks go to Hans Schönemann for resolving various issues encountered with the Singular backend in OSCAR and again to Tommy Hofmann for providing numerous improvements on number-theoretic topics throughout the preparations of this work. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 286237555 – TRR 195.
@@ 2.1. K3 surfaces
Since K3 surfaces are simply connected, linear, algebraic, and numerical equivalence agree and $\mathrm{Pic}(X)=\mathrm{NS}(X)=\mathrm{Num}(X)$. From the exponential sequence, we thus have an embedding $c_{1}\colon\mathrm{Pic}(X)\to H^{2}(X,\mathbb{Z})$ whose image is the subgroup of integral $(1,1)$-classes $H^{1,1}(X)\cap H^{2}(X,\mathbb{Z})$ by Lefschetz’ $(1,1)$-theorem. The smallest primitive sublattice of $H^{2}(X,\mathbb{Z})$ whose complexification contains the period $H^{2,0}(X)$, is called the transcendental lattice $T(X)$. A K3 surface is not necessarily projective, but if it is, then $T(X)=\mathrm{NS}(X)^{\perp}$ in $H^{2}(X,\mathbb{Z})$.
@@
which preserves the intersection form and the Hodge structures, i.e. $\phi_{\mathbb{C}}(H^{2,0}(Y))=H^{2,0}(X)$. The weak Torelli theorem states that $X$ and $Y$ are isomorphic, if and only if $H^{2}(X,\mathbb{Z})$ and $H^{2}(Y,\mathbb{Z})$ are Hodge-isometric.
To formulate the strong Torelli theorem, we need one more definition. The Hodge-isometry $\phi$ is called effective, if it satisfies one of the following three equivalent conditions:
@@
By the Riemann-Roch theorem we have that for $r\in\Delta(X)$ either $r$ or $-r$ is effective. Set $\Delta^{+}(X)=\{r\in\Delta(X)\mid r\mbox{ is effective}\}$. Then $h\in\mathcal{P}_{X}$ is ample if and only if $h.r>0$ for all $r\in\Delta^{+}(X)$. More generally, the connected components of $\mathcal{P}_{X}\setminus\bigcup\{r^{\perp}\otimes\mathbb{R}\mid r\in\Delta(X)\}$ are called Weyl-chambers. The ample cone is one such chamber. The Weyl group $W(X)\subseteq O(\mathrm{NS}(X))$ is generated by the reflections at the hyperplanes $r^{\perp}$. It acts simply-transitively on the set of Weyl-chambers.
@@ 2.2. Enriques surfaces
A complex Enriques surface $Y$ is a compact complex manifold with $h^{1}(Y,\mathcal{O}_{Y})=0$ and $b_{2}(Y)=10$. Its canonical bundle is $\omega_{Y}\neq\mathcal{O}_{Y}$, but $\omega_{Y}^{\otimes 2}\cong\mathcal{O}_{Y}$. Since the canonical bundle is $2$-torsion, it induces a $2:1$ cover $X\to Y$, which turns out to be the universal covering with $X$ a K3 surface. Since $h^{1,1}(Y)=b_{2}(Y)$, it follows from Lefschetz’ Theorem on $(1,1)$-classes and the Kodaira Embedding theorem that all Enriques surfaces are projective.
@@ 3.2. The transcendental lattice
$\{e\in\mathrm{NS}(X)\mid e^{2}=0,h.e=2\}=\{e_{1},e_{2}\}.$
Since the invariant lattice is of rank $1$, $G$ must exchange $e_{1}$ and $e_{2}$ and the stabiliser of $e_{i}$ is of order $G/2=128$. Recall from [shioda] that the trivial lattice $\mathrm{Triv}(e)$ of an elliptic fibration with fiber $e\in\mathrm{NS}(X)$ is the sublattice of $\mathrm{NS}(X)$ spanned by the zero section and fiber components. The Mordell-Weil group is isomorphic to $\mathrm{NS}(X)/\mathrm{Triv}(e)$. To compute the trivial lattice, one chooses any $s\in\mathrm{NS}(X)$ with $s^{2}=-2$ and $s.f=1$. Let $R$ be the root sublattice of $\langle e,s\rangle^{\perp}\subseteq\mathrm{NS}(X)$. Then the trivial lattice is isomorphic to $\mathbb{Z}\langle s,e\rangle\oplus R$. The irreducible root sublattices of $R$ give the ADE types of the singular fibers. For how to compute the fiber types on the Enriques surface $Y$ from the class of $e_{i}$ see [brandhorst-gonzalez, Prop. 5.8]. ∎
@@ 4.1. The first model
It then follows from Proposition 3.2 that $X_{OY}\cong X\cong X_{h}$ and from Proposition 3.6 that $X_{h}/j=:Y$ is isomorphic to the Oguiso-Yu Enriques surface $Y_{OY}$ with an automorphism of minimum entropy.
@@ 4.3. Extracting a Weierstrass model
where $\zeta_{16}$ is a primitive $16$-th root of unity. The associated elliptically fibered K3 surface $X_{e_{1}}$ is defined over $\mathbb{Q}$ and its trivial lattice over $\mathbb{Q}(\zeta_{8})$. However, in order to realize all algebraic cycles for the Néron-Severi lattice $\mathrm{NS}(X_{e_{1}})$, we still need the extension $K/\mathbb{Q}(\zeta_{8})$ of degree two, generated by the algebraic element $\alpha$ with minimal polynomial over $\mathbb{Q}$ given in Equation 6.
Proposition 4.4.
The surface
(9) $\tilde{y}^{2}=s\tilde{x}^{4}+s^{7}-s^{3}$
is birational to the Oguiso-Yu Enriques surface $Y_{OY}$.
Before we go into the proof, we review how to obtain an elliptic fibration of the covering K3 surface greater generality: Let $\widetilde{\pi}\colon Y\to\mathbb{P}^{1}$ an elliptic fibration of an Enriques surface with double fibers over $P_{1},P_{2}\in\mathbb{P}^{1}$, and let $\kappa\colon X\to Y$ be the universal cover. Suppose that $\widetilde{\pi}$ admits a rational bisection. According to [cdl:EnriquesI, Proposition 4.10.10] the Stein-factorization of $\widetilde{\pi}\circ\kappa\colon X\to Y\to\mathbb{P}^{1}$ is given by an elliptic fibration $\pi\colon X_{OY}\to\mathbb{P}^{1}$ followed by a $2:1$ cover $\mathbb{P}^{1}\to\mathbb{P}^{1}$ ramified over $P_{1},P_{2}$.
Proof of Proposition 4.4.
We set $t^{2}=s$, $\tilde{x}=tx_{1}$, $\tilde{y}=t^{3}y_{1}$ in equation (9) and obtain $y_{1}^{2}=x_{1}^{4}+t^{8}-1$. Using the rational point $(x_{1},y_{1})=(1,t^{4})$ we can transform the equation to a short Weierstrass form which one checks to be isomorphic to the one of eq. (8). Putting everything together, we obtain obtain the universal covering morphism $\kappa_{OY}\colon X_{OY}\to Y_{OY}$ by means of the following rational expressions in the Weierstrass coordinates of $X_{OY}$:
$\begin{pmatrix}\tilde{x}\\ \tilde{y}\\ s\end{pmatrix}\mapsto\begin{pmatrix}\frac{-x\cdot t^{9}+x\cdot t+\frac{1}{2}\cdot\sqrt{2}\cdot y\cdot t^{5}-t^{9}+t}{x+\frac{1}{2}\cdot\sqrt{2}\cdot y\cdot t^{4}+t^{8}-1}\\ \frac{-x^{3}\cdot t^{15}-3\cdot x^{2}\cdot t^{15}+3\cdot x^{2}\cdot t^{7}+\frac{1}{2}\cdot y^{2}\cdot t^{15}+\sqrt{2}\cdot y\cdot t^{19}-3\cdot\sqrt{2}\cdot y\cdot t^{11}+2\cdot\sqrt{2}\cdot y\cdot t^{3}+t^{23}-2\cdot t^{15}+t^{7}}{x^{2}+\sqrt{2}\cdot x\cdot y\cdot t^{4}+2\cdot x\cdot t^{8}-2\cdot x+\frac{1}{2}\cdot y^{2}\cdot t^{8}+\sqrt{2}\cdot y\cdot t^{12}-\sqrt{2}\cdot y\cdot t^{4}+t^{16}-2\cdot t^{8}+1}\\ t^{2}\end{pmatrix}$
As ultimate check we confirmed that $\kappa_{OY}=\kappa_{OY}\circ\iota_{OY}$. These computations are carried out in the file
enriques_projection.jl. ∎